# Studying geometry of the ultraluminous X-ray pulsar Swift J0243.6+6124 using X-ray and optical polarimetry

### v1

@@ 3.1 RVM
Previous IXPE data on a number of X-ray pulsars (Doroshenko et al., 2022, 2023; Tsygankov et al., 2022, 2023; Mushtukov et al., 2023) were well described by the rotating vector model (RVM; Radhakrishnan & Cooke, 1969; Meszaros et al., 1988). In this model, the evolution of the PA with pulsar phase is related to the projection of the magnetic dipole on the plane of the sky. If radiation is dominated by the ordinary-mode photons (O-mode), the PA is given by Eq. (30) of Poutanen (2020):
where is the position angle (measured from north to east) of the pulsar angular momentum, is the inclination of the pulsar spin with respect to the line of sight, is the magnetic obliquity, i.e. the angle between the magnetic dipole and the spin axis, and is the phase at which the northern magnetic pole passes in front of the observer (see Fig. 8 for geometry).
If radiation is dominated by the extraordinary mode (X-mode), the PA is rotated by 90° with respect to the O-mode. The general relativistic effects are significant only if the NS rotates at millisecond periods (see Poutanen, 2020; Loktev et al., 2020). The RVM is also theoretically justified, because the polarization vector of the radiation produced at the magnetic poles rotates adjusting to the magnetic field geometry until the photons reach adiabatic radius at a few tens of stellar radii due to the vacuum birefringence (Heyl et al., 2003), where the magnetic field is predominantly dipole (see Sect. 4.2).
@@
We fit the RVM to a given set of Stokes parameters as a function of pulsar phase. The PA distribution does not conform to a normal distribution, hence we use the probability density function from Naghizadeh-Khouei & Clarke (1993):
where is the ‘measured’ PD in units of the error, , and erf is the error function. Here is the prediction of the RVM given by Eq. (1) for a given set of parameters and PA=. The best-fit with the RVM to the pulse-phase dependent is obtained by minimizing the log-likelihood function
with the sum taken over all phase bins for a given observation or summing over all of them. The error on a parameter can be obtained by varying that parameter and fitting all other parameters until is reached. We can also, in principle, use a statistics to evaluate the quality of the fit. This is possible because for highly significant detection of polarization, the PA is distributed almost normally, while the low-significance data points (when the PA distribution is far from normal) contribute very little to the .
First, we fit the RVM to individual observations. The results are given in Table 10. We find that the best-fit parameters are vastly different, e.g., the pulsar inclination for the three observations is 80°, 60°, and 33°. We note that the corresponding values for the best-fits are 25.1, 83.3, and 65.7 for 20 d.o.f. for Obs. 1–3, respectively. Thus, only the fit to Obs. 1 is reasonably good. If, on the other hand, we force the RVM parameters to be the same, we get the best-fit parameters of , , , and (this RVM is shown in Fig. 5e), but an unacceptable fit with . Thus, we find that the RVM does not provide good description of the data.
A similar situation occurred with the brightest transient pulsar observed by IXPE, LS V +44 17/RX J0440.9+4431, which showed the peak luminosity of erg s-1 during its giant outburst in January–February 2023 (Salganik et al., 2023). In this object, significant changes in the RVM parameters were detected (Doroshenko et al., 2023). There, the pulse-phase resolved Stokes and parameters for two epochs of observations traced similar pattern on the plane, but the figures were shifted relative to each other. An additional, phase-independent polarized component was introduced to align the results with the RVM predictions. We now apply the same idea to the J0243 data.
3.2 Two-component polarization model
Similarly to Doroshenko et al. (2023), we assume that there are two polarized components in the J0243 data. The first one is associated with the pulsar and is described by the RVM. The second one is independent of the pulsar phase. We express the absolute Stokes parameters for each observation as a sum of the variable and constant components:
with its PA being .
@@ 3.2 Two-component polarization model
We assume that the pulsar geometry (i.e. RVM parameters) does not change between the observations and the observed changes in the polarization properties are related to the presence of an additional unpulsed polarized component. In order to describe the data from all three observations, in addition to the four RVM parameters, we introduce 6 additional free parameters and , , which describe properties of this constant component. For a given set of and , we can construct the differences , , which are fit by the RVM using maximum likelihood function with the probabilities given by Eq. (2). The best-fit RVM parameters as well as the Stokes parameters of the constant component are given in the top part of Table 11. The quality of the fit is much better than with the RVM model alone. The Akaike information criterion (Akaike, 1974) decreases by , implying that the pure RVM model is less probable that the two-component model.
@@
We find that the PA values of the constant component for the three observations are the same within the errors. This motivates us to perform another fit with the fixed for all three observations (this reduces the number of free parameters by 2). We refitted the data by varying, in addition to the four RVM parameters, three polarized fluxes and the PA . The best-fit parameters are presented in the bottom part of Table 11. We see that the quality of the fit is not getting much worse, with =2.3, i.e. the model is 3 times less probable. The PA of the varying component associated with the pulsar as obtained using Eq. (6) are plotted in Fig. 9 together with the best-fit RVM. In order to obtain the covariance plot for these model parameters we used the affine invariant Markov chain Monte Carlo ensemble sampler emcee package of python (Foreman-Mackey et al., 2013). The resulting posterior distributions are shown in Fig. 10.
@@ 4.1 System geometry and origin of constant polarized component
Our fits demonstrated that polarization cannot be described with a simple RVM. The main reason is that while the PA makes two full turns (by 180°) during one pulsar period during Obs. 3, it shows a double sine wave during Obs. 1 and 2. Such a behavior contradicts the RVM. Furthermore, the quality of the fit is rather bad and the best-fit pulsar geometrical parameters vary strongly between observations (see Table 10).
@@
Following Doroshenko et al. (2023), we assumed that in addition to the direct pulsar radiation described by the RVM, there is a pulse-phase-independent polarized component. The data are much better described by this two-component model with a constant set of RVM parameters. The PA of the constant component in the three observations was found to lie between and , consistent to be the same within the errors. Fixing this PA at the same value gave the best-fit value of . We found that the polarized flux (in units of the average flux) of the constant component is between 1.5% and 3%. Because the contribution of this component to the total flux is unknown, the value of is not well determined. The lower limit on comes from the fact that the flux of the constant component cannot exceed the flux minimum of around 60% of the average flux (see Fig. 1). This results in 3.5–5% depending on the observation, with the value growing inversely proportional to . A lower limit on comes from the condition , which translates to 1.5–3%. If we assume more realistically that (see below), then 5–10%.
@@
The constant component can be associated with scattering in equatorial accretion disk wind (Doroshenko et al., 2023). If the wind half-opening angle is 30°, it occupies half of the sky as seen from the central source. Thus for the Thomson optical depth through the wind of 0.2, 10% of pulsar radiation is scattered by the wind. The PD of the scattered component depends on the disk inclination as (Sunyaev & Titarchuk 1985; Nitindala et al. in prep.). For , the PD is close to 30% and the polarized flux in that case has to be around 3% of the average flux consistent with the data. At lower , a larger contribution of this component to the average flux is required.
@@
The PA of the constant X-ray component is 10° (or ). If its polarization is produced in the accretion disk wind, it is natural to assume that the PA is related to the position angle of the accretion disk axis. On the other hand, polarization in the optical band is likely produced by scattering of the Be-star radiation off the decretion disk (which occupies a much larger solid angle than the accretion disk around a NS) and therefore provides orientation of that disk. Because of the uncertainty in the value of the IS polarization, the intrinsic optical PA is in the range between 20° and 50° (see Tables 9 and 12). The lowest value of (corresponding to the closest star #1 as a proxy for IS polarization) is within from , while the highest PA is clearly inconsistent with that. Thus, there is an indication that the accretion and decretion disks are somewhat misaligned.
@@
From the two-component model fit, we constrained the pulsar inclination at 15°– 40° and the magnetic obliquity to 60°– 80°. The pulsar rotation axis’ position angle is , which is clearly inconsistent with the PA of the constant component . If pulsar radiation escapes in the X-mode, then (or ), which is still far from . Under an assumption that gives the orientation of the accretion disk (and the binary orbit), the difference is related to the misalignment angle between the pulsar spin axis and the orbital axis, which is then about 30°– 50°.
@@
We note that evidence for a misalignment was found also in other pulsars. For example, in LS V +44 17 (Doroshenko et al., 2023), we found and both were inconsistent with implying a misalignment of 75° if the pulsar emits predominantly in the O-mode or at least 15° (for emission in the X-mode). In Her X-1 (Doroshenko et al., 2022; Heyl et al., 2024), the misalignment is at least 25°, but can be 65° or even higher.
@@ 4.2 Polarization – flux anti-correlation
Our fits demonstrate that most of the polarized flux is produced by the pulsating component. Furthermore, the observed polarization has a clear trend of increasing PD with decreasing flux. What would be the most probable reason for such a behavior?
@@
At low luminosities when most of the radiation is produced in two small hotspots at the NS surface close to the magnetic poles, the PD is determined by the structure of the atmosphere and the energy dissipation profile (Doroshenko et al., 2022). The PD is conserved when photons propagate through the magnetosphere and the polarization vector adjusts to the orientation of the magnetic field until the adiabatic radius is reached (e.g., Heyl & Shaviv, 2002; Taverna et al., 2015). The adiabatic radius as a function of photon energy is (Heyl & Caiazzo, 2018)
@@
where is the surface magnetic field strength in units of G and is the NS radius in units of cm. For the surface magnetic field of (Kong et al., 2022) and (e.g., Nättilä et al., 2017; Miller et al., 2020; Annala et al., 2022), it exceeds the NS radius for photons in the IXPE range by more than an order of magnitude. At this distance, the magnetic field is mostly dipole resulting in the PA that follows the RVM, as indeed observed in a number of X-ray pulsars (Doroshenko et al., 2022; Malacaria et al., 2023; Tsygankov et al., 2023; Suleimanov et al., 2023; Mushtukov et al., 2023; Heyl et al., 2024).
@@
At a luminosity close to erg s-1, a number of additional effects come into play. Firstly, the emission region is not point-like anymore, but radiation is produced in an accretion column, which stands above the NS surface. Locally, the polarization direction correlates with the magnetic field direction. However, the PD of the whole column radiation can be smaller that the local values, because of the different energy dissipation profile, varying depth of the vacuum resonance (Gnedin et al., 1978), and resulting mode conversions (Pavlov & Shibanov, 1979; Lai & Ho, 2003; Doroshenko et al., 2022; Lai, 2023). Secondly, a substantial part of the column radiation illuminates the NS surface and is reflected from that (Lyubarskii & Syunyaev, 1988; Poutanen et al., 2013). Because of different relative contributions of the O- and X-modes in the reflected radiation, the total PD is reduced.
And finally, the accretion flow is expected to be optically thick (Mushtukov et al., 2017, 2019) within the NS magnetospheric radius
@@
where is coefficient typically taken to be for the case of disk accretion (Chashkina et al., 2019), is the mass accretion rate in units of g s-1, and is the mass of a NS in solar masses. Under this condition, a substantial fraction of X-ray photons emitted close the NS surface is reprocessed by the optically think envelope created around the NS magnetospheric cavities. Because the size of the magnetosphere exceeds the adiabatic radius, the final polarization of X-ray photons is defined by scattering off the envelope. The PD and PA will depend more on the scattering geometry rather than on the projection of the NS magnetic dipole on the sky. Integrating the Stokes parameters over the envelope, significantly reduces the total PD. Thus in this model, the observed anticorrelation of the PD and flux is a result of variable optical thickness of the envelope: the larger the mass accretion rate and the flux, the larger the optical thickness of envelope and, thus, the larger the fraction of photons experiencing reprocessing by the envelope outside the NS adiabatic radius that reduces the PD.
@@ 5 Summary
Swift J0243.6+6124 was observed by IXPE in July–August of 2023 three times during its outburst. The main results of our study of its polarimetric properties can be summarized as follows:
@@
We showed that evolution of the PA with pulsar phase in Obs. 1 and 2 having double sine-wave structure is inconsistent with the RVM. This brought us to the conclusion that the likely reason for this discrepancy is the presence of the phase-independent polarized component produced, for example, by scattering in the accretion disk wind as was proposed for another bright X-ray pulsar LS V +44 17 (Doroshenko et al., 2023).
@@
Assuming the same RVM parameters for the three observations, we fitted the data with the two-component model and obtained the PA of the constant component between and 15°. Assuming that the PA for all observations is the same, we get . Also assuming that the constant component contributes 10% of the average flux, we get the PD of that component varies between 15% and 30%. We estimated the inclination of the NS rotation axis to the line of sight of 15°– 40°, the magnetic obliquity of 60°– 80°, and the pulsar position angle of .
@@
Using optical polarimetric observations of the source and nearby field stars, we determined the intrinsic PA to lie between 20° and 50°, depending on the choice of a field star to estimate the contribution of the interstellar component. The lowest value of is consistent with within , while the higher PA are clearly inconsistent with that. Associating optical polarization with scattering in the decretion disk, the data indicate a possible misalignment between accretion and decretion disks axes.

### v2

@@ 3.1 RVM
Previous IXPE data on a number of X-ray pulsars (Doroshenko et al., 2022, 2023; Tsygankov et al., 2022, 2023; Mushtukov et al., 2023) were well described by the rotating vector model (RVM; Radhakrishnan & Cooke, 1969; Meszaros et al., 1988). In this model, the evolution of the PA with pulsar phase is related to the projection of the magnetic dipole on the plane of the sky. If radiation is dominated by the ordinary-mode (O-mode) photons, the PA is given by Eq. (30) in Poutanen (2020):
Here is the position angle (measured from north to east) of the pulsar angular momentum, is the inclination of the pulsar spin with respect to the line of sight, is the magnetic obliquity (i.e., the angle between the magnetic dipole and the spin axis), and is the phase at which the northern magnetic pole passes in front of the observer (see Fig. 8 for geometry).
@@
If radiation is dominated by the extraordinary mode (X-mode), the PA is rotated by 90° with respect to the O-mode. The general relativistic effects are significant only if the NS rotates at millisecond periods (see Poutanen, 2020; Loktev et al., 2020). The RVM is also theoretically justified, because the polarization vector of the radiation produced at the magnetic poles rotates adjusting to the magnetic field geometry until photons reach the adiabatic radius at a few tens of stellar radii due to the vacuum birefringence (Heyl et al., 2003), where the magnetic field is predominantly dipole (see Sect. 4.2).
We fit the RVM to a given set of Stokes parameters as a function of pulsar phase. The PA distribution does not conform to a normal distribution, and hence we use the probability density function from Naghizadeh-Khouei & Clarke (1993),
where is the measured PD in units of the error, , and erf is the error function. Here is the prediction of the RVM given by Eq. (2) for a given set of parameters and PA=. The best fit with the RVM to the pulse-phase dependent is obtained by minimizing the log-likelihood function
with the sum taken over all phase bins for a given observation or summing over all of them. The error on a parameter can be obtained by varying that parameter and fitting all other parameters until is reached. We can also, in principle, use a statistic to evaluate the quality of the fit. This is possible because for a highly significant detection of polarization, the PA is distributed almost normally, while the low-significance data points (when the PA distribution is far from normal) contribute very little to the .
First, we fit the RVM to individual observations. The results are given in Table 12. We find that the best-fit parameters are vastly different, for example the pulsar inclination for the three observations is 80°, 60°, and 33°. We note that the corresponding values for the best fits are 25.1, 83.3, and 65.7 for 20 degrees of freedom (d.o.f.) for Obs. 1–3, respectively. Thus, only the fit to Obs. 1 is reasonably good. If, on the other hand, we force the RVM parameters to be the same, we get the best-fit parameters of , , , and (this RVM is shown in Fig. 4f), but an unacceptable fit with . Thus, we find that the RVM does not provide a good description of the data.
A similar situation occurred with the brightest transient pulsar observed by IXPE, LS V +44 17/RX J0440.9+4431, which showed the peak luminosity of erg s-1 during its giant outburst in 2023 January–February (Salganik et al., 2023). In this object, significant changes in the RVM parameters were detected (Doroshenko et al., 2023). There, the pulse-phase resolved Stokes and parameters for two epochs of observations traced similar patterns on the plane, but the figures were shifted relative to each other (see Poutanen et al., 2024). An additional, phase-independent polarized component was introduced to align the results with the RVM predictions. We now apply the same idea to the J0243 data.
@@
3.2 Two-component polarization model
Similarly to Doroshenko et al. (2023), we assume that there are two polarized components in the J0243 data. The first is associated with the pulsar and is described by the RVM. The second component is independent of the pulsar phase. We express the absolute Stokes parameters for each observation as a sum of the variable and constant components:
with its PA being .
We assume that the pulsar geometry (i.e., RVM parameters) does not change between the observations and the observed changes in the polarization properties are related to the presence of an additional unpulsed polarized component. In order to describe the data from all three observations, in addition to the four RVM parameters, we introduce six additional free parameters and , , which describe properties of this constant component. For a given set of and , we can construct the differences , , which are fit by the RVM using maximum likelihood function with the probabilities given by Eq. (3). The best-fit RVM parameters as well as the Stokes parameters of the constant component are given in the top part of Table 13. The quality of the fit is much better than with the RVM model alone. The Akaike information criterion (AIC; Akaike, 1974) decreases by , implying that the pure RVM model is less probable that the two-component model.
@@ 3.2 Two-component polarization model
We found that the PA values of the constant component for the three observations are the same within the errors. This motivated us to perform another fit with the fixed for all three observations (which reduced the number of free parameters by 2). We refitted the data by varying, in addition to the four RVM parameters, three polarized fluxes and the PA . The best-fit parameters are presented in the bottom part of Table 13. We see that the quality of the fit is not getting much worse, with =2.3, which means that the model is three times less probable. The PA of the varying component associated with the pulsar as obtained using Eq. (7) are plotted in Fig. 9 together with the best-fit RVM. In order to obtain the covariance plot for these model parameters we used the affine invariant Markov chain Monte Carlo ensemble sampler emcee package of python (Foreman-Mackey et al., 2013). The resulting posterior distributions are shown in Fig. 10.
@@ 4.1 System geometry and origin of constant polarized component
Our fits demonstrated that polarization cannot be described with a simple RVM. The main reason is that while the PA makes two full turns (by 180°) during one pulsar period during Obs. 3, it shows a double sine wave during Obs. 1 and 2. This behavior contradicts the RVM. Furthermore, the quality of the fit is rather bad and the best-fit pulsar geometrical parameters vary strongly between observations (see Table 12).
@@
Following Doroshenko et al. (2023), in addition to the direct pulsar radiation described by the RVM, we assumed that there is a pulse-phase-independent polarized component. The data are much better described by this two-component model with a constant set of RVM parameters. The PA of the constant component in the three observations was found to lie between and , consistent with being the same within the errors. Fixing this PA at the same value gave the best-fit value of . We found that the polarized flux (in units of the average flux) of the constant component is between 1.5% and 3%. Because the contribution of this component to the total flux is unknown, the value of is not well determined. The lower limit on comes from the fact that the flux of the constant component cannot exceed the flux minimum of around 60% of the average flux (see Fig. 9). This results in 3.5–5% depending on the observation, with the value growing inversely proportional to . A lower limit on comes from the condition , which translates to 1.5–3%. If we assume more realistically that (see below), then 5–10%.
@@
The constant component can be associated with scattering in the equatorial accretion disk wind (Doroshenko et al., 2023). If the wind half-opening angle is 30°, it occupies half of the sky as seen from the central source. Thus, for the Thomson optical depth through the wind of 0.2, about 10% of pulsar radiation is scattered by the wind. The PD of the scattered component depends on the disk inclination as (Sunyaev & Titarchuk 1985; Nitindala et al. in prep.). For , the PD is close to 30% and the polarized flux in that case has to be around 3% of the average flux consistent with the data. At lower , a larger contribution of this component to the average flux is required.
@@
The PA of the constant X-ray component is 10° (or ). If its polarization is produced in the accretion disk wind, it is natural to assume that the PA is related to the position angle of the accretion disk axis. On the other hand, polarization in the optical band is likely produced by scattering of the Be-star radiation off the decretion disk (which occupies a much larger solid angle than the accretion disk around a NS), and therefore provides the orientation of that disk. Because of the uncertainty in the value of the IS polarization, the intrinsic optical PA is in the range between 20° and 50° (see Tables 11 and 14). The lowest value of (corresponding to the closest star #1 as a proxy for IS polarization) is within of , while the highest PA is clearly inconsistent with that. Thus, there is an indication that the accretion and decretion disks are somewhat misaligned.
@@
From the two-component model fit, we constrained the pulsar inclination at 15°– 40° and the magnetic obliquity to 60°– 80°. The pulsar rotation axis position angle is , which is clearly inconsistent with the PA of the constant component . If pulsar radiation escapes in the X-mode, then (or ), which is still far from . Under the assumption that gives the orientation of the accretion disk (and the binary orbit), the difference is related to the misalignment angle between the pulsar spin axis and the orbital axis, which is then about 30°– 50°.
@@
We note that evidence for a misalignment was also found in other pulsars. For example, in LS V +44 17 (Doroshenko et al., 2023) we found , and both were inconsistent with implying a misalignment of 75° if the pulsar emits predominantly in the O-mode or at least 15° for emission in the X-mode. In Her X-1 (Doroshenko et al., 2022; Heyl et al., 2024), the misalignment is at least 25°, but can be 65° or even higher.
@@ 4.2 Polarization–flux anti-correlation
Our fits demonstrate that most of the polarized flux is produced by the pulsating component. Furthermore, the observed polarization has a clear trend of increasing PD with decreasing flux. We wondered what the most probable reason for such a behavior could be.
@@
At low luminosities when most of the radiation is produced in two small hotspots at the NS surface close to the magnetic poles, the PD is determined by the structure of the atmosphere and the energy dissipation profile (Doroshenko et al., 2022). The PD is conserved when photons propagate through the magnetosphere and the polarization vector adjusts to the orientation of the magnetic field until the adiabatic radius is reached (e.g., Heyl & Shaviv, 2002; Taverna et al., 2015). The adiabatic radius as a function of photon energy is (Heyl & Caiazzo, 2018; Taverna & Turolla, 2024)
@@
where is the surface magnetic field strength in units of G, and is the NS radius in units of cm. For the surface magnetic field of (Kong et al., 2022) and (e.g., Nättilä et al., 2017; Miller et al., 2020; Annala et al., 2022), it exceeds the NS radius for photons in the IXPE range by more than an order of magnitude. At this distance, the magnetic field is mostly dipole resulting in the PA that follows the RVM, as indeed observed in a number of X-ray pulsars (Doroshenko et al., 2022; Malacaria et al., 2023; Tsygankov et al., 2023; Suleimanov et al., 2023; Mushtukov et al., 2023; Heyl et al., 2024).
@@
At a luminosity close to erg s-1, a number of additional effects come into play. First, the emission region is not point-like anymore, but radiation is produced in an accretion column, which stands above the NS surface. Locally, the polarization direction correlates with the magnetic field direction. However, the PD of the whole column radiation can be smaller that the local values, because of the different energy dissipation profile, varying depth of the vacuum resonance (Gnedin et al., 1978), and resulting mode conversions (Pavlov & Shibanov, 1979; Lai & Ho, 2003; Doroshenko et al., 2022; Lai, 2023). Second, a substantial part of the column radiation illuminates the NS surface and is reflected from that (Lyubarskii & Syunyaev, 1988; Poutanen et al., 2013). Because of the different relative contributions of the O- and X-modes in the reflected radiation, the total PD is reduced. Finally, the accretion flow is expected to be optically thick (Mushtukov et al., 2017, 2019) within the NS magnetospheric radius
@@
where is coefficient typically taken to be for the case of disk accretion (Chashkina et al., 2019), is the mass accretion rate in units of g s-1, and is the mass of a NS in solar masses. Under this condition, a substantial fraction of X-ray photons emitted close the NS surface is reprocessed by the optically thick envelope created around the NS magnetospheric cavities. Because the size of the magnetosphere exceeds the adiabatic radius, the final polarization of X-ray photons is defined by scattering off the envelope. The PD and PA depend more on the scattering geometry than on the projection of the NS magnetic dipole on the sky. Integrating the Stokes parameters over the envelope significantly reduces the total PD. Thus, in this model the observed anticorrelation of the PD and flux is a result of variable optical thickness of the envelope: the higher the mass accretion rate and the flux, the larger the optical thickness of envelope, and thus the larger the fraction of photons experiencing reprocessing by the envelope outside the NS adiabatic radius that reduces the PD.
@@ 5 Summary
Swift J0243.6+6124 was observed by IXPE in 2023 July–August three times during its outburst. The main results of our study of its polarimetric properties can be summarized as follows:
@@
We showed that evolution of the PA with pulsar phase in Obs. 1 and 2 having a double sine wave structure is inconsistent with the RVM. This brought us to the conclusion that the likely reason for this discrepancy is the presence of the phase-independent polarized component produced, for example by scattering in the accretion disk wind, as was proposed for another bright X-ray pulsar LS V +44 17 (Doroshenko et al., 2023; Poutanen et al., 2024).
@@
Assuming the same RVM parameters for the three observations, we fitted the data with the two-component model and obtained the PA of the constant component between and 15°. Assuming that the PA for all observations is the same, we find . Also assuming that the constant component contributes 10% of the average flux, we find that the PD of that component varies between 15% and 30%. We estimated the inclination of the NS rotation axis to the line of sight of 15°– 40°, the magnetic obliquity of 60°– 80°, and the pulsar position angle of .
@@
Using optical polarimetric observations of the source and nearby field stars, we determined that the intrinsic PA lies between 20° and 50°, depending on the choice of field star to estimate the contribution of the interstellar component. The lowest value of is consistent with within , while the higher PA are clearly inconsistent with that value. Associating optical polarization with scattering in the decretion disk, the data indicate a possible misalignment between accretion and decretion disks axes.